If f is continuous except at finitely many points in I,then f is Riemann-integrable.
1
Expert's answer
2020-10-28T18:51:17-0400


Let "\isin" >0. Put M=sup"_{x\in[a,b]}" f

.Now choose a partition so that the total length of the intervals containing the discontnuities of f

f is smaller than "\dfrac{\in}{2M}"


The contribution from the intervals with discontinuities of f

f is smaller than "\dfrac{\in}{2}"

Since f is continuous on the rest, it is fairly easy to engineer an argument that supplies a partition whose upper and lower sums differ by less than "\dfrac{\in}{2}"


Assemble the pieces and you have Riemann integrability.

so f is Riemann integral.